# Recoverable strain from reverse plasticity

### v1

@@ V.2 Strain recovery after stress switch-off
If the strain were held constant, the macroscopic effect of these plastic yielding events from negative to zero $l$ would be a positive growth in stress, via Eqn. 9, as elements hop from negative to zero local stress. Conversely, under the conditions of constant stress that pertain after stress switch off, $\sigma=0$ and $\dot{\sigma}=0$, a corresponding negative strain rate is required to advect elements in the negative direction in their traps, in order to counterbalance this positive plastic stress change. Accordingly, as weight is lost from the negative $l$ part of the distribution in moving from the magenta to green curves, these curves also simultaneously advect leftwards. We therefore now understand the strain recovery that arises after stress switch-off within the SGR model to stem from these ‘reverse local plastic yielding events’.
With this picture of the basic physics of strain recovery via reverse plasticity in mind, we shall now explore the dependence of the phenomenon on each of the three control parameters in our study: the amplitude of imposed stress $\sigma_{0}$, the degree to which the sample plastically strains forward while the stress is held imposed $\Delta\gamma_{\rm f}$, and the age of the sample $t_{\rm w}$ before stress switch-on. We shall consider first in Fig. 5 the total strain $\Delta\gamma_{\rm rec}$ recovered plastically at long times $t-t_{\rm stop}\to\infty$, before examining the time-dependence of the recovery process in Fig. 6. Recall that $\Delta\gamma_{\rm f}$ and $\Delta\gamma_{\rm rec}$ are indicated on the left and hand sides respectively of Fig. 3b): a value $\Delta\gamma_{\rm rec}=0$ corresponds to zero plastic strain recovery (beyond the initial elastic recoil $\Delta\gamma=-\gamma_{0}$ at the instant of switch-off), whereas a value $\Delta\gamma_{\rm rec}=\Delta\gamma_{\rm f}$ would indicate full recovery.
@@
To understand the trends observed in Fig. 5 we recall that, immediately prior to stress switch-off, any elements that remain unyielded since the stress was imposed have local strain $l=\sigma_{0}+\Delta\gamma_{\rm f}$, as in the rightmost lobe of the cyan distribution in Fig. 4. (This actually has a slightly smeared distribution of strain values of small width $l_{0}$. As noted above, this is simply for the purposes of being able to plot the distribution graphically and does not change the basic physics.) We shall call these “group I” elements. In contrast, those elements that yielded at least once during the initial forward creep (“group II”) have local strain values in the range $0<l<\Delta\gamma_{\rm f}$, giving the leftmost lobe of that distribution. At the instant of stress switch-off, the whole distribution of shifts to the left, with each $l\to l-\sigma_{0}$. Immediately after switch-off, therefore, the unyielded group I elements have positive local strain $l=\Delta\gamma_{\rm f}$, as in the (slightly smeared) rightmost lobe of the magenta distribution in Fig. 4. They furthermore reside in traps of a depth set by the age $t_{\rm w}$, and with a rate of yielding $r\sim 1/t_{\rm w}$. In contrast, the yielded group II elements have local strain values in the range $-\sigma_{0}<l<\Delta\gamma_{\rm f}-\sigma_{0}$ immediately post switch-off, as in leftmost lobe of the magenta distribution in Fig. 4. And in being freshly yielded, they reside in shallow traps with a rate of yielding $r=O(1)$. The relative proportion of elements in group II compared to those in group I increases with increasing forward strain $\Delta\gamma_{\rm f}$.
@@
A different trend is however observed for larger values of the scaled forward strain prior to switch-off, $\Delta\gamma_{\rm f}/\sigma_{0}\gtrapprox 1$. In this case, for large imposed stresses, the ultimate plastic strain recovery is actually observed to be negative: the system strains plastically further forwards as a function of time $t>0$ since switch-off. The overall strain recovery is therefore non-monotonic, with the initially backward elastic recoil at the instant of switch-off $t=t_{\rm stop}$ followed by this forward plastic straining for $t>t_{\rm stop}$, reminiscent of observations of non-monotonic stress relaxation after straining Hendricks et al. (2019); Sudreau et al. (2022); Murphy et al. (2020); Mandal et al. (2021). (The total recovery is still however negative, with the initial elastic recoil always exceeding any subsequent forward plastic straining, for all parameter regimes we have explored.)
@@
VI Conclusions
@@
We have reported the counter-intuitive phenomena that a significant component of the strain recovery post switch-off arises not via the recoil of elasticity, but instead by reverse plasticity. We have also showed that the time-evolution of strain post switch-off is not always monotonic, but can be non-monotonic (albeit in relatively rare parameter regimes): the plastic straining that takes place for times after switch-off $t>t_{\rm stop}$ does not always occur in the negative direction, as in the initial elastic recoil at $t=t_{\rm stop}$, but can sometimes in fact continue to accumulate in the same forward direction as the originally imposed stress. This provides a potential counterpart, in this imposed-stress protocol, to recent observations of non-monotonic stress relaxation after the switch off of an imposed strain Hendricks et al. (2019); Sudreau et al. (2022); Murphy et al. (2020); Mandal et al. (2021).
@@
We have discussed these observations in terms of the evolution of the SGR model’s population of elastoplastic elements. In particular, we have shown that strain recovery arises via elements that locally yielded during the initial forward plastic creep, resetting their local strains from positive values to zero as they did so. These elements then subsequently yield again post switch-off, resetting their local strains (and stresses) from negative values to zero. It is this plastic yielding from from negative to zero stress values that leads to global strain recovery, in order to maintain the global stress of the sample as a whole equal to zero post switch-off.
The physics just described is strongly reminiscent of the concept of reversible plasticity, which has been studied extensively in the concept of oscillatory shear: localised plastic T1 events that arise in the forward direction during the positive part of a strain cycle exactly reverse during the reverse part of the cycle  Galloway et al. (2020); Keim and Arratia (2014, 2013); Lundberg et al. (2008, 2008); Regev et al. (2013); Priezjev (2016); Szulc et al. (2022); Nagasawa et al. (2019); Khirallah et al. (2021). It should be noted, however, that the SGR model as studied here does not provide any spatial resolution of elemental positions. Accordingly, we are not able to make any statement as to whether a plastic event at a specific location later exactly reverses at the same location. Instead, we make only the weaker statement that some elements that yield during the initial forward creep later do so in reverse during recovery, recognising that our mean field approach does not track the spatial location of the elements. For this reason, we have chosen to use the weaker term reverse plasticity rather than the the term reversible plasticity. Future work should study creep and recovery in spatially resolved models, to establish a stronger link to reversible plasticity, with its precise connotations of elemental positions.
@@
Acknowledgements — The authors thank Mike Cates and Gareth McKinley for helpful comments on an early draft of the manuscript. We also thank Andrew Clarke for discussions and SLB (Schlumberger Cambridge Research Ltd.) for support. This project has received funding from the European Research Council (ERC) under the European Union’s Horizon 2020 research and innovation programme (grant agreement No. 885146).

### v2

@@ V.2 Strain recovery after stress switch-off
If the strain were held constant, the macroscopic effect of these plastic yielding events from negative to zero $l$ would be a positive growth in stress, via Eqn. 9, as elements hop from negative to zero local stress. Conversely, under the conditions of constant stress that pertain after stress switch off, $\sigma=0$ and $\dot{\sigma}=0$, a corresponding negative strain rate is required to advect elements in the negative direction in their traps, in order to counterbalance this positive plastic stress change. Accordingly, as weight is lost from the negative $l$ part of the distribution in moving from the magenta to green curves, these curves also simultaneously advect leftwards. We therefore now understand the slow ongoing strain recovery that arises after the initial elastic recoil post switch-off to stem from the reverse local plastic yielding of elements left with negative stress values by that initial recoil. We further recognise these elements to be those that had previously yielded plastically in the forward direction while the material was under load. In this way, counter-intuitively, stress relaxation that occurs plastically while a material is under load can still in fact contribute to strain recovery after the load is removed.
With this basic physics in mind, we shall now explore the dependence of the phenomenon on each of the three control parameters in our study: the amplitude of imposed stress $\sigma_{0}$, the degree to which the sample plastically strains forward while the stress is held imposed $\Delta\gamma_{\rm f}$, and the age of the sample $t_{\rm w}$ before stress switch-on. We shall consider first in Fig. 5 the total strain $\Delta\gamma_{\rm rec}$ recovered plastically at long times $t-t_{\rm stop}\to\infty$, before examining the time-dependence of the recovery process in Fig. 6. Recall that $\Delta\gamma_{\rm f}$ and $\Delta\gamma_{\rm rec}$ are indicated on the left and hand sides respectively of Fig. 3b): a value $\Delta\gamma_{\rm rec}=0$ corresponds to zero plastic strain recovery (beyond the initial elastic recoil $\Delta\gamma=-\gamma_{0}$ at the instant of switch-off), whereas a value $\Delta\gamma_{\rm rec}=\Delta\gamma_{\rm f}$ would indicate full recovery.
@@
To understand the trends observed in Fig. 5 we recall that, immediately prior to stress switch-off, any elements that remain unyielded since the stress was imposed have local strain $l=\sigma_{0}+\Delta\gamma_{\rm f}$, as in the rightmost lobe of the cyan distribution in Fig. 4. (This actually has a slightly smeared distribution of strain values of small width $l_{0}$. As noted above, this is simply for the purposes of being able to plot the distribution graphically and does not change the basic physics.) We shall call these “group I” elements. In contrast, those elements that yielded at least once during the initial forward creep (“group II”) have local strain values in the range $0<l<\Delta\gamma_{\rm f}$, giving the leftmost lobe of that distribution. At the instant of stress switch-off, the whole distribution of shifts to the left, with each $l\to l-\sigma_{0}$, corresponding to a fast elastic recoil. Immediately after switch-off, therefore, the unyielded group I elements have positive local strain $l=\Delta\gamma_{\rm f}$, as in the (slightly smeared) rightmost lobe of the magenta distribution in Fig. 4. They furthermore reside in traps of a depth set by the age $t_{\rm w}$, and with a rate of yielding $r\sim 1/t_{\rm w}$. In contrast, the yielded group II elements have local strain values in the range $-\sigma_{0}<l<\Delta\gamma_{\rm f}-\sigma_{0}$ immediately post switch-off, as in leftmost lobe of the magenta distribution in Fig. 4. And in being freshly yielded, they reside in shallow traps with a rate of yielding $r=O(1)$. The relative proportion of elements in group II compared to those in group I increases with increasing forward strain $\Delta\gamma_{\rm f}$.
@@
A different trend is however observed for larger values of the scaled forward strain prior to switch-off, $\Delta\gamma_{\rm f}/\sigma_{0}\gtrapprox 1$. In this case, for large imposed stresses, the ultimate plastic strain recovery is actually observed to be negative: the system strains plastically further forwards as a function of time $t>0$ since switch-off. The overall strain recovery is therefore non-monotonic, with the initially backward elastic recoil at the instant of switch-off $t=t_{\rm stop}$ followed by this forward plastic straining for $t>t_{\rm stop}$, reminiscent of observations of non-monotonic stress relaxation after straining Hendricks et al. (2019); Sudreau et al. (2022); Murphy et al. (2020); Mandal et al. (2021). The total recovery is still however negative, with the initial elastic recoil always exceeding any subsequent forward plastic straining, for all parameter regimes we have explored.
@@
Figure 7: Strain response with non-zero local strain frustration, for parameter values otherwise as in Fig. 3. a) Effect of non-zero width $l_{p}$ to the distribution of strains selected by any element immediately post-hop, with $l_{p}=0.0,0.1\cdots 1.0$ in curves downwards at the right. b) Effect of a non-zero width $l_{0}$ to the distribution of strains in the sample at time $t=0$, after sample preparation and before shearing commences, with $l_{0}=0.0,0.1\cdots 1.0$ in curves upward at the right.
So far, we have assumed the distribution of local strains present in the sample at time $t=0$, before shearing commences, to be a delta function of zero mean. (An exception can be found in Fig. 3, where we assumed a narrow Gaussian simply to facilitate plotting.) We have likewise assumed the distribution of strains selected by any element immediately post-hop to be a delta function of zero mean. We now consider the more general case in which the distribution before shearing commences is a Gaussian of zero mean and standard deviation $l_{0}$, and the post-hop distribution is a Gaussian of zero mean and standard deviation $l_{p}$. (Our results so far therefore have $l_{0}=l_{p}=0$.) The parameter $l_{0}$ thus now represents the degree of local strain frustration present in the sample before shearing starts, arising from whatever complicated ageing or annealing dynamics took places during sample preparation. Smaller values of $l_{0}$ correspond to better ordered, more strongly annealed samples. Likewise, $l_{p}$ represents the frustration experienced by elements in attempting to relax their local strains during local rearrangement events.
In Fig. 7, we explore the effect of non-zero values of $l_{0}$ and $l_{p}$ on the basic phenomenon of recoverable strain arising from reverse plastic events. Shown are results counterpart to those in Fig. 3, but now for a range of $l_{p}$ values in panel a) and $l_{0}$ in  3b). As can be seen, the basic phenomenon is quite robust to varying these quantities. Indeed, the degree of strain recovery from reverse plastic events increases with increasing post-hop frustration $l_{p}$, as might be expected intuitively: the left hand lobe of the turquoise distribution (for example) in Fig. 4a) will now be more smeared, including some weight to the left of the origin. When elastically shifted leftward to form the magenta distribution in b), more weight will be present at more negative $l$ values, resulting in faster yielding of elements with more strongly negative $l$, and more negative global straining (i.e. more strain recovery) to compensate. And although the effect of plastic strain recovery decreases with increasing initial frustration $l_{0}$ in panel Fig. 7b), it nonetheless persists to values of this parameter $O(1)$.
VI Interplay of initial (visco)elastic recoil and slow ongoing plasticity
So far, we have used the SGR model in its original form, considering only the elastoplastic stress arising from the population of SGR elements. In this version of the model, the initial response to an applied step stress is instantaneous: elements shift elastically within their traps at the instant of the step. Accordingly, in Fig. 3 (say) we saw an instantaneous elastic recoil at the time the stress is switched off, followed by the later slow recovery arising from reverse plastic yielding of elements left with negative local strand by the initial recoil. (The instantaneous elastic upward step during switch-on at $t=0$ is not seen on the logarithmic time axis in Fig. 3.)
In practice, however, drag against a background solvent will render the timescale for this first part of the strain response non-zero, such that the initial recoil is now viscoelastic rather than purely elastic. An obvious question is then to what extent the initial (visco)elastic recoil and later reverse plastic yielding of elements with negative local stresses during strain recovery after stress switch-off overlap in time, or whether they occur truly sequentially, with the fast (visco)elastic recoil simply acting to leave elements with negative local stresses that later relax to give reverse plastic events. To answer this question, we now consider the total stress $\Sigma=\sigma+\eta\dot{\gamma}$ to comprise the usual elastoplastic stress $\sigma$ as predicted by the SGR model in Eqn. 5, plus a Newtonian solvent contribution of viscosity $\eta$.
Figure 8: Strain response with non-zero background solvent viscosity $\eta$, for parameter values otherwise as in Fig. 3. a) Forward strain creep while the stress is held imposed. b) Strain recovery after the stress is switched off. c) Number of plastic events per element accumulated while the stress is held imposed. (d) Number of plastic events per element accumulated after the stress is switched off. Viscosity values $\eta=10^{n}$ with $n=-3.0,-2.5\cdots 0.0$ in curves left to right in a), with the same colour code across all sub-panels.
Fig. 8 shows results for forward creep and reverse strain recovery with model parameter values otherwise as in Fig. 3, but now with a non-zero solvent viscosity $\eta$. As can be seen, this indeed imposes a finite timescale for the initial rise in strain after stress switch-on (panel a), and for the initial (now visco)elastic strain recoil after stress switch-off (panel b). The timescale for each of these processes scales as $\eta/k$, tending to zero in the limit $\eta\to 0$. It is in this limit that the initial elastic recoil and later reverse plastic events become truly distinct: for finite $\eta$, the two processes merge somewhat.
Panels c) and d) show the number of plastic events $n$ accumulated per element on average during forward creep and reverse recovery. Consistent with plastic events being thermally activated (as well as strain-induced) in the SGR model (rather than purely strain-induced, as would be the case in an athermal system), these signals show only modest dependence on the strain (and so on $\eta$). As can be seen, the plastic part of the forward creep and – importantly to the key message of this work – of the reverse recovery, arises in each case only once $n$ approaches $O(1)$, denoting significant plastic yielding.
We have seen, then, that the initial (visco)elastic recoil post-switch off and the later slow plastic strain recovery will only be truly time-separated in the limit of zero solvent viscosity $\eta\to 0$. In either case (for $\eta$ finite or zero), however, it is clear that these two stages interact, in the sense that the initial (visco)elastic recoil creates the state in which some elements have negative local strain values. It thereby creates the conditions out of which the slower plastic part of the strain recovery can later emerge.
VII Implications for constitutive modelling
Finally, we consider for the purpose of pedagogical comparison the response of a simplified fluidity model to the same stress protocol as studied for the SGR model in this work. The model is defined as follows. The total stress comprises an elastoplastic contribution $\sigma$ and a Newtonian solvent contribution of viscosity $\eta$:
$\Sigma=\sigma+\eta\dot{\gamma}.$ (10)
The elastoplastic stress evolves with dynamics
$\frac{d\sigma}{dt}=G\dot{\gamma}-\frac{\sigma}{\tau},$ (11)
in which the relaxation timescale $\tau$ has its own dynamics
$\frac{d\tau}{dt}=1-|\dot{\gamma}|(\tau-\tau_{0}).$ (12)
Here $G$ is a constant elastic modulus and $\tau_{0}$ a constant microscopic timescale. (Artificially setting the relaxation time to infinity then gives Kelvin-Voigt dynamics. Instead artificially setting the viscosity to zero and the relaxation time to a constant gives a Maxwell model.) The dynamics assumed for the relaxation timescale $\tau$ confers a yield stress, as in SGR, with a flow curve given by $\sigma=G(1+\dot{\gamma}\tau_{0})$. In the absence of any imposed flow the model displays ageing in which the stress relaxation timescale $\tau$ increases linearly in time, also as in SGR.
Figure 9: Strain response of simplified fluidity model with non-zero background solvent viscosity $\eta$. a) Forward strain creep while the stress is held imposed. b) Strain recovery after the stress is switched off. c) Integrated rate of plasticity accumulated while the stress is held imposed. (d) Integrated rate of plasticity accumulated after the stress is switched off. Viscosity values $\eta=10^{n}$ with $n=-3.0,-2.5\cdots 0.0$ in curves left to right in a), with the same colour code across all sub-panels. $\Sigma=1.4$, $t_{\rm w}=10^{3}$, $\Delta\gamma_{\rm f}=1.4$.
[…]
This has important implications for constitutive modeling, in showing that the phenomenon studied in this work can be captured only by models that evolve a full distribution of stress values (or multiple moments of such a distribution), rather than a single sample-averaged stress.
VIII Conclusions
@@
We have reported the counter-intuitive phenomena that a significant component of the strain recovery post switch-off arises, after an initially fast elastic recoil, via the slow reverse plastic yielding of elements left with negative local stresses by that recoil. We have also showed that the time-evolution of strain post switch-off is not always monotonic, but can be non-monotonic (albeit in relatively rare parameter regimes): the plastic straining that takes place for times after switch-off $t>t_{\rm stop}$ does not always occur in the negative direction, as in the initial elastic recoil at $t=t_{\rm stop}$, but can sometimes in fact continue to accumulate in the same forward direction as the originally imposed stress. This provides a potential counterpart, in this imposed-stress protocol, to recent observations of non-monotonic stress relaxation after the switch off of an imposed strain Hendricks et al. (2019); Sudreau et al. (2022); Murphy et al. (2020); Mandal et al. (2021).
@@
We have discussed these observations in terms of the evolution of the SGR model’s population of elastoplastic elements. In particular, we have shown that plastic strain recovery (after the initial fast elastic recoil) arises via elements that locally yielded during the initial forward plastic creep, resetting their local strains from positive values to zero as they did so. These elements then acquire negative local stresses during the fast elastic recoil, and later yield again post switch-off, resetting these local stresses from negative values to zero. It is this plastic yielding from from negative to zero stress values that leads to global plastic strain recovery, in order to maintain the global stress of the sample as a whole equal to zero post switch-off. In this way, elements that yield plastically in the forward direction while a material is under load can still contribute to strain recovery after load removal.
The phenomenon reported here has important implications for constitutive modelling, in showing that a full distribution of stress values (or multiple moments of such a distribution) is needed to capture it. Any model that evolves just a single average stress will fail to describe it.
The physics described in this work is reminiscent of but distinct from the concept of reversible plasticity, which has gained increasing attention in the recent physics literature. Indeed, a key concept in the rheology of amorphous materials is that of intermittent local plastic events, sometimes called T1 events, arising at soft spots or shear transformation zones Falk and Langer (2011), coupled by elastic stress propagation Picard et al. (2004). Within this framework, oscillatory shear experiments on jammed particle rafts identified three different regimes of imposed strain amplitude Galloway et al. (2020); Keim and Arratia (2014, 2013). At low amplitudes, the material responds elastically, with no plastic events. At high amplitudes, irreversible plastic events occur within each cycle, with significant cycle-to-cycle stroboscopic changes in particle positions. More surprisingly, at intermediate strain amplitudes these experiments report reversible local plastic events within each cycle. Such events dissipate energy, consistent with the notion of plasticity, but exactly reverse in the opposite part of the cycle, with zero stroboscopic cycle-to-cycle change in particle positions. Reversible plastic events have also been reported in experiments on foams Lundberg et al. (2008), in particle simulations Lundberg et al. (2008); Regev et al. (2013); Priezjev (2016); Szulc et al. (2022); Nagasawa et al. (2019) and in elastoplastic models Khirallah et al. (2021). Reversible plastic limit cycles have likewise been studied in models of hysterons Keim and Paulsen (2021) and iterated maps Mungan and Witten (2019). For a review, see Ref. Reichhardt et al. (2022).
It should be noted, however, that the mean field SGR model as studied here does not provide any spatial resolution of elemental positions. Accordingly, we are not able to make any statement as to whether a plastic event at a specific location later exactly reverses at the same location. Furthermore, because plastic events in SGR are thermally activated as well as being strain-induced, exact reversibility under straining is unlikely, except in the athermal limit $x\to 0$ – in which limit, however, predictions of sustained slow creep are difficult to obtain.
Therefore, we do not claim a close correspondence between the plasticity that arises in this work in the reverse direction during strain recovery post-switch off and the truly reversible plasticity in which individual elements exactly retrace their trajectories, as seen under oscillatory shear in athermal systems. Instead, we make only the weaker statement that some elements that yield during the initial forward creep later do so in reverse during recovery. For this reason, we have been careful to use the weaker term reverse plastic events rather than the the term reversible plasticity, as coined in the literature concerning oscillatory shear in athermal systems. Indeed, we use the term “reverse” in this work to mean “in the opposite direction”, as distinct from “reversible” meaning “an event capable of being reversed”. Future work should study creep and recovery in spatially resolved athermal models of amorphous materials, to establish whether there is a stronger link to reversible plasticity, with its precise connotations of exact localised reversibility.
@@
Acknowledgements — The authors thank Mike Cates, Andrew Clarke and Gareth McKinley for comments on an early draft of the manuscript, and for helpful discussions. We also thank SLB (Schlumberger Cambridge Research Ltd.) for support. This project has received funding from the European Research Council (ERC) under the European Union’s Horizon 2020 research and innovation programme (grant agreement No. 885146). We thank the anonymous reviewers for their insightful comments that improved the clarity of our descriptions.
Author declarations — The authors have no conflicts to disclose.